# Bi-Fact: A Bidirectional Factorization-based Evaluation of Intent Extraction from UI Trajectories

### v2

@@ Abstract
Evaluating intent extraction from GUIs demands accurate, fine-grained metrics. This paper introduces Bi-Fact, a novel method that decomposes intents into atomic facts and performs bidirectional comparisons to assess precision and recall. Experiments demonstrate Bi-Fact’s superior correlation with human judgments compared to existing metrics, establishing a more robust evaluation framework for UI-driven intent understanding.
@@
Sapir Caduri sapir@google.com
@@ 1 Introduction
However, these applications depend on the accuracy of individual extracted facts. For instance, consider the following gold intent "book an business flight to Paris" if the model prediction was "book a flight to Paris", although its not fully correct most of the facts were correctly identified. Knowing them allows the model to generate followup actions for the user and learn useful memory items about the users activity. Therefore, a robust evaluation framework is needed to assess the accuracy of intent predictions at a granular, fact-based level.
Recent research has begun to address intent extraction from UI interactions Berkovitch et al. (2024); Sun et al. (2024); Huang et al. (2024); Martínez-Rojas et al. (2024); Zhang et al. (2024); Yang et al. (2024). However, evaluating these extracted intents remains a hurdle. Existing metrics struggle to capture the nuances of UI-driven intents, especially regarding factual accuracy.
Traditional text similarity measures like BLEU Papineni et al. (2002) and ROUGE Lin (2004) rely on lexical overlap and are ill-suited for capturing semantic equivalence in UI contexts. Similarly, semantic similarity metrics and inference-based methods like Natural Language Inference (NLI) Bowman et al. (2015) often fall short. While full-intent level NLI can assess overall entailment between two intents, it may not capture the fine-grained factual discrepancies crucial for downstream tasks. For example, a detailed request for a trip booking on Monday and a similar request for that trip on Tuesday are related but do not strictly entail or contradict each other, potentially leading NLI models to miss the pivotal differences in dates. Fact-level evaluation methods, such as FactScore Min et al. (2023), offer a more granular approach by decomposing generated text into factual claims and checking if these are supported by a reference text.
However, evaluating intent understanding in UI interactions requires considering both the recall and precision of factual information. Consider a UI trajectory where a user searches for "restaurants near Central Park." A model extracts the intent "find restaurants open after 10 PM in Central Park." While assessing whether this extracted intent recalls the facts from the user’s search is valuable, it doesn’t address whether the extracted intent introduces overly specific information. A comprehensive evaluation requires assessing both precision (are all extracted facts correct and relevant?) and recall (are all user-intended facts captured?).
@@
In this work, we propose Bi-Fact, a novel fact-level evaluation method specifically designed for assessing the equivalence of intents extracted from UI interactions. Bi-Fact leverages a fact-based comparison, enabling a more detailed and accurate evaluation than existing metrics. The core idea involves decomposing intents into their constituent facts and comparing these facts’ validity. We demonstrate that Bi-Fact exhibits significantly high correlation with human judgments of fact-level intent equivalence. This indicates that Bi-Fact provides a more robust and informative evaluation framework for accurately assessing intent predictions in UI interaction scenarios.
@@
Figure 1: Example of Bi-Fact Evaluation Process: Comparing a Gold (reference) intent with a Predicted (model-generated) intent. Each intent is decomposed into atomic facts. Facts are compared bidirectionally comparison represented with checkmarks (✓) for derived and crosses (✗) for missed facts. Recall, precision, and F1 score can be computed from the bidirectional comparison.
@@ 2 The Bi-Fact method of Evaluating Intent Extraction
Evaluating the accuracy of intent extraction requires a nuanced, fact-based approach.
@@ 2.1 Approach
At a high-level, given a gold intent and a predicted intent, we would like to decompose both intents into atomic facts, representing single, indivisible pieces of information. For example, the intent "book a flight to Paris for a weekend business trip" might be decomposed into facts like "destination is Paris," "trip type is business," and "duration is weekend." We then perform a bidirectional evaluation: checking whether each fact in the annotation is derived from the gold intent, and conversely, whether each fact in the gold intent can be derived from the annotation, considering synonyms, paraphrases, and implicit information. The flow is summarized in Figure 1.
@@ 2.2 Automatic Evaluation with Bi-Fact
We introduce Bi-Fact, an automatic evaluation method utilizing a large language model (LLM). The methodology consists of the following two primary stages, with the assessment stage further divided into three reasoning steps:
Part 1: Factoring the Gold Sentences: Initially, the gold sentences are factorized into atomic facts. During this stage, the factual content of the gold sentences is frozen, providing a stable basis for subsequent comparison of future models. The prompt used for this factorization is detailed in the Appendix  2.
Part 2: Factual Coverage Assessment: This stage involves assessing the factual coverage of the predicted sentences against the factorized gold sentences. This assessment is conducted in three distinct reasoning steps:
@@
    Reasoning Step 1: Factoring the Prediction: The LLM first decomposes the predicted intent into its atomic facts, ensuring a consistent level of granularity for comparison.
@@
    Reasoning Step 2: Assessing the Expert Facts: The LLM analyzes each expert-provided fact (from the gold reference) and labels it as either ’derived’ or ’not derived’ based on its presence or absence in the factorized prediction. This step focuses on recall, measuring whether all facts from the gold standard are captured.
@@
    Reasoning Step 3: Assessing the Predicted Facts: The LLM analyzes each fact in the factorized prediction and labels it as either ’derived’ or ’not derived’ based on its presence or absence in the gold standard. This step focuses on precision, ensuring that the predicted intent does not contain extraneous or incorrect information.
The prompt used for this process is detailed in Appendix  3.
Bi-Fact’s strength lies in its atomic-level operation. This granular approach breaks down complex sentences for accurate comparison, preserves meaning by focusing on factual content, and provides fine-grained assessments with ’derived’/’not derived’ labels. By reducing ambiguity and estimating recall/precision through atomic-level evaluation, Bi-Fact ensures a comprehensive assessment.
3 Evaluating Effectiveness of Bi-Fact
@@ 3.1 Datasets
We utilize two datasets to evaluate the effectiveness of our proposed Bi-Fact metric.
@@
To evaluate our proposed metric, we utilize the dataset from Berkovitch et al. (2024), which provides manually annotated intent data. The dataset comprises 100 examples sampled from the test sets of Mind2Web Deng et al. (2024) and AitW Rawles et al. (2024). Each example includes a human-generated intent describing the UI trajectory and model-generated intent predictions from Gemini and GPT. This results in a total of 400 intent pairs. Each pair was manually annotated to indicate whether both intents achieve the same result. We allocated 10% of this data for development to calibrate the thresholds of the different comparison methods, and the remaining 90% was used as a test set. This dataset facilitates a comprehensive evaluation of intent comparison across manual annotations from multiple datasets and model-generated predictions from various sources, making it suitable for assessing the performance of our proposed metric.
@@
To validate the correctness of our automatic method, we conducted manual factorization on a small dataset of 36 intent pairs. We manually compared intents and calculated F1 scores based on these annotations. In Section 3.3, we demonstrate a high correlation between our automatic method and the manual annotations.
3.2 Comparison to other approaches on Intent-match
We compare Bi-Fact against standard text similarity metrics, including BLEU Papineni et al. (2002), ROUGE Lin (2004), METEOR banerjee2005meteor, T5 cosine similarity (ni2021sentence), and NLI-based methods Bowman et al. (2015). We binarized the continuous scores using thresholds that yielded the best performance in terms of F1 score on the development set (testing 30 equally spaced increments between 0.01 and 1.0). For the NLI-based method, we used a bidirectional approach, taking the average of the NLI scores for both directions (premise vs. hypothesis and hypothesis vs. premise). We also include the Automatic Rater (AutoRater) introduced in Berkovitch et al. (2024). To assess the effectiveness of these metrics, we evaluate the correlation between each of these scores and human judgments on the intent-match data using recall, precision, F1 score, and kappa. The results are in Table 1. Bi-Fact achieves the highest F1 score (0.722) and Kappa coefficient (0.508) among all tested metrics, indicating better agreement with human judgments. The results highlight the limitations of purely lexical overlap based metrics for capturing the nuances of intent in UI interactions.
@@ 3.2 Comparison to other approaches on Intent-match
Table 1: Comparing different text similarity metrics’ effectiveness at evaluating intent match. Bidirectional comparisons are used, and continuous scores are binarized using a threshold.
@@ 3.3 Comparison using Fact-level annotations
We further validated Bi-Fact by analyzing its correlation with manual fact-level annotations. This analysis was conducted on the Manual Fact-Level dataset described in Section 3.1, providing insights into Bi-Fact’s ability to capture fine-grained factual information that is highly crucial to the downstream tasks. Bi-Fact’s F1 score demonstrated a strong correlation with these human fact-level annotations, achieving a Pearson correlation of 0.781 (p < 0.001). This substantially outperformed the second-best method, NLI, which reached only 0.517 (p = 0.001).

### v3

@@ Abstract
Evaluating user intent extraction from GUI interactions demands accurate, fine-grained metrics. This paper introduces Bi-Fact, a novel method that decomposes intents into atomic facts and performs bidirectional comparisons with reference intents in order to assess precision and recall. Experiments demonstrate Bi-Fact’s superior correlation with human judgments compared to existing metrics, establishing a more robust automated evaluation framework for UI-driven intent understanding.
@@
Sapir Caduri 1  Anatoly Efros 1  Noam Kahlon 1  Danielle Cohen 1 Yoni Halpern 1Ido Dagan 1,2 1Google Research  2Bar-Ilan University {sapir,talef,kahlonn,daniellecn,yhalpern,idodagan}@google.com
@@ 1 Introduction
Notably, these applications depend on the accuracy of individual extracted facts. For instance, consider the following gold intent “book a business-class flight to Paris” where the model prediction was “book a flight to Paris”. Although the prediction is not fully correct, some of the facts within the gold intent were correctly identified. Obtaining them would allow downstream models, for example, to generate follow-up actions for the user and learn useful personal profile items relating to the users activity. Therefore, a robust evaluation framework is needed to assess the accuracy of intent predictions at a granular, fact-based level.
Recent research has begun to address intent extraction from UI interactions, developing various models for the task Berkovitch et al. (2024); Sun et al. (2024); Huang et al. (2024); Martínez-Rojas et al. (2024); Zhang et al. (2024); Yang et al. (2024). In this task setting, a model receives as input a trajectory of a GUI session, consisting of screenshots and user actions, and generates a predicted natural-language intent description, which is supposed to fully and precisely describe the user intents which were fulfilled during the session. However, evaluating these extracted intents remains a hurdle. Existing metrics struggle to capture the nuances of UI-driven intents, especially regarding fine-grained factual accuracy and completeness. Traditional text similarity measures like BLEU Papineni et al. (2002) and ROUGE Lin (2004) rely on lexical overlap and are ill-suited for capturing semantic equivalence in UI contexts. On the other hand, sentence-level semantic similarity metrics and inference-based methods like classifying textual-entailment with Natural Language Inference (NLI) models Dagan et al. (2006); Bowman et al. (2015) also fall short. While full-intent level NLI can assess overall entailment between two intents, it does not capture the fine-grained factual discrepancies crucial for downstream tasks.
For example, a detailed request for a trip booking on Monday and a similar request for that trip on Tuesday are related but do not strictly entail or contradict each other, potentially leading NLI models to miss the pivotal differences in dates. Fact-level evaluation methods, such as Pyramid Nenkova and Passonneau (2004) and its automated variants Gao et al. (2019); Yang et al. (2016); Harnly et al. (2005),, or FactScore Min et al. (2023), offer a more granular approach by comparing the generated and reference texts at a granular level of factual claims, which are extracted from the respective text. However, while Pyramid is a recall-oriented measure and FactScore is precision-oriented, evaluating intent understanding requires considering both the recall (are all user-intended facts captured?) and precision (are all extracted facts correct and relevant?) of the extracted factual information relative to the gold intent.
@@
In this work, we propose Bi-Fact, a novel fact-level evaluation method specifically designed for assessing intents extracted from UI interactions, enabling a more detailed and accurate evaluation than prior metrics utilized for this setting. The core idea is to first decompose both predicted and gold intents into their constituent facts, and then to test for the support for each fact in the counterpart intent. We demonstrate that Bi-Fact exhibits significantly high correlation with human judgments of fact-level intent equivalence, providing a robust and informative evaluation framework for assessing intent extraction in UI interaction scenarios.
@@
Figure 1: Example of Bi-Fact Evaluation Process: Comparing a Gold (reference) intent with a Predicted (model-generated) intent. Each intent is decomposed into atomic facts. Facts are compared bidirectionally for their support by the other text, as represented with checkmarks (✓) for implied facts and crosses (✗) for missed facts. Recall, precision, and F1 scores are computed from the bidirectional comparison.
@@ 2.1 Approach
At a high-level, given a gold intent and a predicted intent, we would like to decompose both intents into atomic facts, representing single, indivisible pieces of information. For example, the intent “book a flight to Paris for a weekend business trip”, might be decomposed into facts like “book a flight”, “destination is Paris”, “trip type is business”, and “duration is weekend”. We then perform a bidirectional evaluation: checking whether each fact in the predicted intent is implied (entailed) from the gold intent (precision), and conversely, whether each fact in the gold intent can is implied from the predicted intent (recall). The computation flow is summarized in Figure 1.
@@ 2.2 Automatic Evaluation with Bi-Fact
We automate the Bi-Fact evaluation protocol utilizing a large language model (LLM). The methodology consists of the following two primary parts, with the assessment step further divided into three reasoning stages:
Part 1: Factorizing the Gold Intents: Initially, given a dataset with gold intents (along with their corresponding trajectories), the gold intents are factorized into atomic facts. Following this stage, the factual decomposition of the gold intents is fixed, providing a stable basis for subsequent comparative evaluations of the predictions made by various models. The prompt used for this factorization is detailed in Appendix A (Figure  2).
Part 2: Factual Coverage Assessment: This part is performed when evaluating the predicted intents generated by an intent extraction model. It consists of three stages, first decomposing the predicted intent into a set of individual facts, and then assessing the recall and precision of the predicted intent relative to the gold intent. This phase is performed in a single prompt (Figure 3).
@@
    Stage 1: Factorizing the predicted intent: The LLM first decomposes the predicted intent into its atomic facts. As an additional input, the LLM receives the decomposition of the gold intent (as fixed in Part 1), and is instructed to decompose the predicted intent in accordance with the given decomposition of the gold intent, to increase their consistency for comparison.
@@
    Stage 2: Recall assessment: The LLM examines each gold fact (from the decomposed gold intent, obtained in Part 1), and labels it as either implied (entailed) or not-implied from the predicted intent, where the latter is considered as a whole. Recall is then measured as the proportion of implied facts among the gold facts.
@@
    Stage 3: Assessing the predicted facts: The LLM examines each predicted fact (from the decomposed predicted intent, obtained in Stage 1), and labels it as either implied (entailed) or not-implied from the gold intent, where the latter is considered as a whole. Precision is then measured as the proportion of implied facts among the predicted facts.
Bi-Fact’s strength lies in its atomic-level operation. This granular approach breaks down complex sentences for fact-level comparison, thus supporting a comprehensive assessment of both precision and recall, while providing precise information regarding which facts are supported and which are not, in both the gold and predicted intents.
3 Evaluating Bi-Fact
@@ 3.1 Datasets
We utilize two datasets to evaluate the effectiveness of our proposed Bi-Fact metric. These datasets include human assessments of predicted intents, in comparison to given gold intents. Thus, these datasets allow measuring the agreement between these human intent evaluations and automatic evaluation measures, which are applied to the same gold and predicted intent pairs.
@@
To evaluate our proposed metric, we utilize the dataset from Berkovitch et al. (2024), which provides manually annotated intent data. The dataset comprises 100 examples sampled from the test sets of Mind2Web Deng et al. (2024) and AitW Rawles et al. (2024). Each example includes a human-generated intent describing the UI trajectory and model-generated intent predictions from Gemini and GPT. This results in a total of 400 gold-predicted intent pairs. Each pair was manually annotated with a binary judgement, of whether both intents indicate the same task, in the sense that two trajectories, fulfilling the gold and predicted intent, respectively, would achieve the same outcome. Accordingly, this dataset is suitable to assess automatic evaluation measures that yield a binary judgement of whether a predicted intent, as a whole, is semantically equivalent to the gold intent, with respect to the task it defines. We allocated 10% of this data as development set to calibrate the thresholds of the different intent evaluation methods, and the remaining 90% were used as a test set. This dataset facilitates a comprehensive evaluation of intent comparison across manual annotations from multiple datasets and model-generated predictions from various sources, making it suitable for assessing the performance of our proposed intent evaluation metric.
@@
To validate the correctness of our automatic method at the granular fact-level assessment, we conducted a thorough manual factorization on a small dataset of 36 intent pairs. We then manually assessed the entailment of the decomposed intent facts across the gold and predicted intents in each instance, yielding the respective precision and recall scores and the corresponding F1 scores based on these annotations. In Section 3.3, we demonstrate a high correlation between our automatic method and these manual annotations.
3.2 Comparison to other evaluation approaches on binary Intent-match
For the binary evaluation judgment (whether a predicted intent is equivalent to a gold intent, as a whole), we compare Bi-Fact against standard text similarity metrics, including BLEU Papineni et al. (2002), ROUGE Lin (2004), METEOR Banerjee and Lavie (2005), T5 cosine similarity (Ni et al., 2021), and NLI-based methods Honovich et al. (2022). We binarized the continuous scores using thresholds that yielded the best performance in terms of F1 score on the development set (testing 30 equally spaced increments between 0.01 and 1.0). For the NLI-based method, we used a bidirectional approach, taking the average of the NLI scores for both directions (premise vs. hypothesis and hypothesis vs. premise). We also include the Automatic Rater (AutoRater) introduced in Berkovitch et al. (2024). For methods that yield a numeric score (rather than a binary judgement), we tune a hyper-parameter of a score threshold for positive judgement (match vs. non-match) over the development set of the Intent-Match data. To assess the effectiveness of these metrics, we evaluate the correlation between each of these scores and human judgments on the intent-match data using recall, precision, F1 score, and kappa. The results are in Table 1. Bi-Fact notably achieves the highest F1 score (0.722) and Kappa coefficient (0.508) among all tested metrics, indicating better agreement with human judgments. The results highlight the limitations of purely lexical overlap based metrics for capturing the nuances of intent in UI interactions.
@@ 3.2 Comparison to other evaluation approaches on binary Intent-match
Table 1: Comparing different text comparison metrics’ effectiveness at evaluating binary intent match. Bidirectional comparisons are used, and continuous scores are binarized using a threshold.
@@ 3.3 Comparison using Fact-level annotations
We further validated Bi-Fact by analyzing its correlation with manual fact-level annotations. This analysis was conducted on the Manual Fact-Level dataset described in Section 3.1, providing insights into Bi-Fact’s ability to capture fine-grained factual information, which is highly crucial and informative for downstream tasks, as discussed in the Introduction. Bi-Fact’s F1 score demonstrates a strong correlation with the human fact-level annotations, achieving a Pearson correlation of 0.781 (p < 0.001). This substantially outperformed the second-best method, NLI, which reached only 0.517 (p = 0.001).
@@ 4 Discussion
Appendix A Prompts
Figure 2: Example prompt for the fact decomposition task
Figure 3: Assessment prompt part 1 - 3 step instructions
Figure 4: Assessment prompt part 2 - Output Structure
Figure 5: Assessment prompt part 3 - An Example (one-shot)